Paul’s Perfect Peugeot says they’ll sell you a brand new Italian “Iron Man” motor scooter for $1,699. Financing is available, and the terms are 10% down and payments of $46.57 a month for 40 months. What interest rate is Paul charging you?

(please use excel formulas)
Answer #1

loan amount=1699*(1-10%)=1529.10

What interest rate is Paul charging you using excel function

=RATE(nper,pmt,pv,fv)

=(RATE(40,46.57,-1529.10,0))*12

=12.00%

the above is answer..